# Real-to-Sim via End-to-End Differentiable Simulation and Rendering

### v1

@@ Abstract
Identifying predictive world models for robots in novel environments from sparse online observations is essential for robot task planning and execution in novel environments. However, existing methods that leverage differentiable simulators to identify world models are incapable of jointly optimizing the shape, appearance, and physical properties of the scene. In this work, we introduce a novel object representation that allows the joint identification of these properties. Our method employs a novel differentiable point-based object representation coupled with a grid-based appearance field, which allows differentiable object collision detection and rendering. Combined with a differentiable physical simulator, we achieve end-to-end optimization of world models, given the sparse visual and tactile observations of a physical motion sequence. Through a series of benchmarking system identification tasks in simulated and real environments, we show that our method can learn both simulation- and rendering-ready world models from only a few partial observations.
@@ I Introduction
Building a predictive world model of how a robot’s actions can affect the surrounding environment is essential for planning and control. However, constructing such a model from raw observations in novel real-world environments remains challenging. The shapes, appearances, and physical parameters of the objects in the scene are usually partially observable, and robots are typically limited in time and computational resources.
@@
Recently, there has been growing interest in learning world models from large offline datasets of action-labeled videos using generative modeling techniques [1, 2, 3, 4]. However, these models are susceptible to distribution shifts and cannot infer latent properties such as the coefficient of friction. In addition, they are not physically consistent and cannot provide information such as contact forces, which are essential for downstream tasks. Meanwhile, system identification of world models aims to leverage strong priors to estimate the parameters of world models from noisy, sparse, and partial observations. However, identifying general-purpose world models requires the understanding of geometry, appearance, and physical properties of the environment, such as object-segmentation labels, geometric shapes, mass-inertial properties, and frictional coefficients.
@@
Many existing works take advantage of differentiable simulators as strong priors that allow efficient system identification of physical parameters [5, 6, 7]. Differentiable simulators allow the gradient-based optimization of mass-inertial properties and frictional coefficients, in order for a physical motion sequence to match sparse robot observations. However, these works assume known shapes and appearances of the objects involved and do not allow the algorithm to adapt the shape, appearance, and physical properties simultaneously.
@@
On the other hand, we have witnessed recent advances in learnable appearance models, such as Neural Radiance Fields (NeRF) [8] and Gaussian Splatting (GS) [9]. These methods build the rendering equation into a learnable representation to enable the identification of shapes and appearances with general geometries and topologies, from raw observations. However, rigid body simulators [10] typically require the use of volumetric representations such as convex hulls to detect collisions, while appearance models such as neural fields and Gaussian splats are incompatible with these conventional representations. In summary, to the best of the authors’ knowledge, no existing method allows the simultaneous identification of physical, geometric, and appearance properties of a scene from sparse robot observations.
@@
Figure 1: Our method jointly optimizes the shape, appearance, and physical parameters of a world model in the form of a rigid body simulator, from the visual and tactile observations of a single robot push.
@@
To tackle these challenges, this work presents an object representation that is compatible with general-purpose rigid body simulators and allows the joint optimization of geometry, appearance, and physical parameters. Our proposed representation is the combination of a recently proposed point-based shape representation Shape-as-Points (SaP) [11] and a grid-based appearance property field. SaP parameterizes an object’s geometry and topology using a set of surface points along with normal directions. It then uses differentiable Poisson reconstruction to recover a smooth indicator field, which can be converted to a mesh using a differentiable marching cubes algorithm. The texture of the vertices of the mesh is then obtained by interpolating the appearance grid. Employing the mesh in a differentiable rigid body simulator [12] that provides gradients for the physical parameters and contact points of the objects, our method constructs a fully differentiable pipeline for jointly optimizing the geometry, appearance, and physical parameters. To summarize, our contributions are:
@@
  • A jointly differentiable representation of objects’ shape, appearance, and physical properties.
@@
We evaluate our method on a series of benchmarking problems in both simulated and real-world environments. As illustrated in Fig. 1, the results show that our method can infer faithful world models from a single episode of robot interactions with the environment.
@@ II-A Differentiable Rigid Body Simulator
Rigid body simulators are essential tools in robotics and engineering for testing and verification, perception, control, and planning. Traditional rigid body simulators are not differentiable, but recently, there have been many attempts at building differentiable rigid body simulators [12, 13, 14, 15, 16, 17] for facilitating downstream system identification, robot planning, and policy optimization tasks. Different strategies are adopted to enable the calculation of gradients for the underlying non-differentiable contact dynamics, including employing a smooth contact model [17, 14, 15], using sub-gradients of the linear complementarity problem [13], and implicit gradients of nonlinear optimization [12, 16]. The majority of these methods, however, do not provide gradients with respect to the geometry, with the exception of  [16, 12, 14]. In this work, we adopt the simulator proposed by Strecke et al. [12] for its physical realism, numerical stability, and fast computation from GPU acceleration.
@@ II-B Learnable Geometry and Appearance Models
Learning 3D geometry and appearance models from 2D raw images is vital to robots’ understanding of the physical world. Earlier research has focused on learning only the 3D geometries without appearance, including point-cloud-based [18] and convex-hull-based [19] models. Neural radiance fields (NeRF) is the first method that enables a continuously learnable model for the full 3D appearance of objects and scenes, which uses neural networks to parameterize the spatial appearance properties and implicitly learn the 3D geometry. More recently, Kerbl proposed Gaussian splitting (GS) [9], a non-parametric method that represents the appearance of the scene with 3D Gaussians and significantly improves the training and rendering speeds due to their fit for fast GPU/CUDA-based rasterization. These algorithms are capable of learning detailed 3D object and scene appearances from sparse image-based observations. However, NeRF represents objects with a neural field and lacks clear definition of rigid object surfaces, and the case is similar with GS. Therefore, while there are some initial attempts at integrating them with rigid body simulators [20, 21], research for robust and physically correct collision detections with these models is still ongoing. In addition, NeRF and GS require multiple views of the image observations for supervision, which is unrealistic in typical robotics applications.
Our method is based on the shape representation SaP [11], with a set of sampled points with normals to represent the object surface. It then reconstructs the indicator field by Poisson reconstruction. A mesh can then be reconstructed from the indicator field using a differentiable marching cubes algorithm. Coupled with a 3D appearance property grid, the mesh vertex textures can be assigned. Compared to standard 3D representations such as point clouds, which do not allow volumetric collision detection, or meshes, which do not allow topology changes during optimization, our SaP-based methods enjoy the best of both worlds. Combined with a differentiable renderer, our object representation then achieves end-to-end image-based shape optimization.
@@ II-C World Models
In this work, we interpret the term world models broadly and regard traditional system identification methods [22] as world modeling as well, where only the dynamics are identified from full state information. However, to support diverse downstream robot tasks in the real world, world models need to be built from raw observations and support both accurate dynamics prediction and photorealistic novel view synthesis. As previously mentioned, while existing works have identified world models from raw image observations using differentiable simulators, none supports simultaneous optimization of geometry, appearance, and physical parameters. Another line of works closely related to ours is image-based generative world modeling. These works aim to predict the next RGB frame based on the current frame and action, crafted by training on diverse datasets with generative modeling techniques such as variational autoencoders [1, 2] and diffusion models [3, 4]. The key differences between our method and these works are that our simulation, grounded in physics, is always physically consistent and is a general-purpose simulator that can also provide information such as contact forces. On the other hand, purely data-driven world models generalize poorly to novel scenarios and have limited capabilities for downstream robot tasks.
@@
Figure 2: Overview of the proposed fully differentiable pipeline for world model identification from sparse robot observations. Our object representation couples an oriented point cloud and a 3D appearance properties grid. Through a differentiable Poisson solver and differentiable marching cubes, the oriented point cloud is converted to an indicator grid and then a mesh, whose vertex textures are interpolated from the appearance properties grid. Feeding the object mesh, physical parameters, the terrain point cloud, and the robot pushing action into a differentiable rigid body simulator and renderer, the predicted scenes can be rendered. Calculating the loss against observed RGB-D images, the scene shape, appearance, and physical parameters are jointly optimized with gradient descent.
@@ III Problem Definition
In this section, we describe our formulation of world model identification, or real-to-sim. We assume the environment consists of a rigid object and a rigid terrain, whose physical properties and appearances are parameterized by $\theta$. A robot, equipped with joint encoders and end-effector force sensors, interacts with the object at $T$ sparse time instances: $t_{1},\cdots,t_{T}$, with a fixed time step $\delta t$. At each time step, the robot observes its end-effector pose $e^{t}\in\text{SE}(3)$ and contact force $f^{t}\in\mathbb{R}^{3}$. Further, the robot is equipped with an RGB-D camera with known intrinsics that observes the object through image $o^{t}\in\mathbb{R}^{H\times W\times 4}$ at camera pose $c^{t}\in\text{SE}(3)$. We further assume an image segmentation mask $m^{t}\in\mathbb{R}^{H\times W\times 4}$ is provided for the object, robot, and terrain. Therefore, the robot observations are a sequence $\mathcal{O}=\{\langle t,e^{t},f^{t},o^{t},c^{t},m^{t}\rangle\}$, and our goal is to estimate $\theta$ from the set of sparse observations $\mathcal{O}$. We formulate this problem as a physics-constrained optimization by introducing a full-fledged physics simulator function $q^{i+1}=g(q^{i},u^{i},\theta)$ that can differentiate through objects’ appearance, geometry, and physical parameters. Here $q^{i}\in\text{SE}(3)\times\text{SE}(3)$ is the object and robot end-effector poses at timestep $t$ and $u^{i}$ is the applied robot force to the object, which is equal in magnitude to the sensed contact force but opposite in direction.
@@ IV-A Differentiable Object Representation
We find that the SaP framework [11], when augmented by additional appearance properties poses an ideal representation for our purpose. Briefly, this framework represents the object using a point cloud with normal on the object surface, denoted as $\mathcal{P}=\{\langle p\in\mathbb{R}^{3},n\in\mathbb{R}^{3}\rangle\}$. These normal directions induce a discrete vector field $v(x)=\sum_{\langle p,n\rangle\in\mathcal{P}}n\mathbb{I}[x=p]$. SaP then uses Poisson reconstruction [23] to recover an underlying implicit indicator field $\chi(x)$ that describes the solid geometry and matches its gradient field with $v(x)$ by solving the variational problem:
@@
which amounts to solving the Poisson equation $\Delta\chi=\nabla\dot{v}$. SaP discretizes the indicator field $\chi$ on a uniform grid domain $\Omega$, which allows the efficition solution of $\chi$ via GPU-accelerated Fast Fourier Transform (FFT) with well-defined derivatives.
The indicator field $\chi$ is then transformed to a mesh $\mathcal{M}$ with a differentiable marching cubes algorithm [24]. Collision detection can then be easily achieved with standard techniques for meshes. To enable appearance modeling, we further augment with an appearance properties grid, with the same grid resolution as the one storing the indicator field $\chi$. The appearance property is then propagated to the mesh vertices via tri-linear interpolation. In this work, we only store and render the color field, denoted as $\psi$, but other appearance properties can be incorporated in the same manner as required by more advanced differentiable rendering equations. The mesh can then be rendered using any differentiable renderer framework such as [25, 26], for which we use the open source implementation in PyTorch3D [27]. Specifically, at the time instance $t$, we invoke the renderer with the object transformed to $q^{t}$ and the camera transformed to $c^{t}$. In summary, our parameterization of the objects’ physics and appearance is defined as:
@@
where the first two parameters are mass-inertial properties and frictional coefficients, defining the physical properties, and the last two parameters are the oriented point cloud and color field, defining the objects’ appearance.
@@ IV-B Differentiable Simulator
For our application, we only consider unconstrained rigid body dynamics with dry frictional contacts. Note that additional constraints such as soft bodies and articulated objects can be readily incorporated into our framework and its differentiation has been well-studied, e.g. in [15].
@@
with $M(q^{i})$ being the generalized mass matrix, $C(q^{i},\dot{q}^{i})$ being the centrifugal, Coriolis, and gravitational force, $J^{i},J_{\perp},J_{\parallel}$ being the Jacobian matrix for the external, normal, and tangent contact forces at all the detected contact points, respectively. Finally, $\tau_{\perp},\tau_{\parallel}$ are the contact forces. At each time step, a mixed LCP problem is solved to calculate the constraint forces $\tau_{\perp},\tau_{\parallel}$, yielding the final acceleration $\ddot{o}^{i}$, and we then integrate the configuration forward in time [30, 31] as:
@@
$\begin{cases}0\leq\tau_{\perp}&\perp J_{\perp}^{T}\dot{q}^{i+1}\geq 0\\ 0\leq\tau_{\parallel}&\perp\lambda e+J_{\parallel}^{T}\dot{q}^{i+1}\geq 0\\ 0\leq\lambda&\perp\mu\tau_{\perp}-e^{T}\tau_{\parallel}\geq 0,\end{cases}$ (4)
with $e$ being the unit vector and $\mu$ being the frictional coefficient. $\lambda$ is an auxiliary variable deciding the static and slide frictional state. To differentiate through the simulator, we adopt the differentiation technique proposed by [13, 12], where the result of the LCP is made differentiable by solving with a primal-dual method and performing sensitivity analysis at the solution to yield derivatives with respect to the problem data. In this way, the derivatives propagates the gradient information to the Jacobian matrix $J_{\perp,_{\parallel}}$, and finally to the object geometric data $\langle\mathcal{P},\chi\rangle$. In summary, Eqn. 2,3,4 defines our differentiable simulator function $g$. In particular, we adopt the differentiable simulator proposed by Strecke et al. [12] for its fast implementation on GPU.

### v3

# One-Shot Real-to-Sim via End-to-End Differentiable Simulation and Rendering

@@ Abstract
Identifying predictive world models for robots in novel environments from sparse online observations is essential for robot task planning and execution in novel environments. However, existing methods that leverage differentiable programming to identify world models are incapable of jointly optimizing the geometry, appearance, and physical properties of the scene. In this work, we introduce a novel rigid object representation that allows the joint identification of these properties. Our method employs a novel differentiable point-based geometry representation coupled with a grid-based appearance field, which allows differentiable object collision detection and rendering. Combined with a differentiable physical simulator, we achieve end-to-end optimization of world models, given the sparse visual and tactile observations of a physical motion sequence. Through a series of world model identification tasks in simulated and real environments, we show that our method can learn both simulation- and rendering-ready world models from only one robot action sequence.
@@ I Introduction
An accurate internal model of a robot about how its actions can affect the surrounding environment is essential for robot planning and control. Such a model, which we refer to as a world model, needs to render realistic raw observations such as RGB images from arbitrary viewpoints and predict consistent and accurate physical interactions. However, constructing such a model from raw observations in novel real-world environments remains challenging as it requires the identification of the geometry parameters that describe the shape of all objects (e.g. vertices and faces of a mesh), appearance parameters that define how the objects look when rendered (e.g. color and reflectance), and physical parameters (e.g. mass) of the objects in the scene. These parameters are usually partially observable, and robots are typically limited in time and computational resources.
@@
Recently, there has been growing interest in learning world models from large offline datasets of action-labeled videos using generative modeling techniques [1, 2, 3, 4]. However, these black-box models are susceptible to distribution shifts and cannot infer properties such as the coefficient of friction. In addition, they are not physically consistent and cannot provide physical information such as contact forces, which are essential for downstream tasks. Meanwhile, an alternative approach that identifies the geometry, appearance, and physical parameters (GAP) of the environment with strong priors coming from knowledge of physics can result in a generalizable and physically consistent world model.
@@
Many existing works take advantage of differentiable simulators as strong physics priors that allow efficient identification of physical parameters such as inertia and coefficient of friction [5, 6, 7]. Differentiable simulators allow the gradient-based optimization of mass-inertial properties and frictional coefficients to match a physical motion sequence to sparse robot observations. However, these works assume known geometries and appearances of the objects in the scene and do not allow the algorithm to adapt the GAP simultaneously.
@@
On the other hand, we have witnessed recent advances in learnable geometry and appearance models, such as Neural Radiance Fields (NeRF) [8] and Gaussian Splatting (GS) [9]. These methods build the rendering equation into a learnable representation to enable the identification of geometries and appearances from raw observations. However, rigid body simulators [10] typically require the use of volumetric representations with a clear definition of object surfaces such as convex hulls to detect collisions and penetration depths. Unfortunately, NeRF and GS are incompatible with the requirements of rigid body simulators since NeRF represents objects with a continuous neural field and GS with individual 3D Gaussians. To the best of the authors’ knowledge, no existing method allows the simultaneous identification of the GAP properties of a world model of rigid objects from sparse robot observations.
@@
Figure 1: From the visual and tactile observations of a single robot push (top), our method jointly optimizes the shape, appearance, and physical parameters of a world model consisting of rigid objects in the form of a rigid body simulator (bottom, the robot arm is not rendered in this picture and the end-effector is treated as a floating blue sphere robot).
@@
To address these challenges, this work presents an object representation that is compatible with general-purpose rigid body simulators and allows the joint optimization of GAP. As shown in Fig. 1, based on this representation, our work enables the identification of a rigid world model in the form of a full-fledged rigid body simulator from the observations of one robot push. Our proposed representation is the combination of a recently proposed point-based shape representation Shape-as-Points (SaP) [11] and a grid-based appearance field. SaP parameterizes an object’s geometry and topology using a set of surface points along with normal directions. It then uses differentiable Poisson reconstruction to recover a smooth indicator field of object occupancy, which can be converted to a mesh using a differentiable marching cubes algorithm [12]. The texture of the vertices of the mesh is then obtained by interpolating the appearance grid. Employing the mesh in a differentiable rigid body simulator [13] that provides gradients for the physical parameters and contact points of the objects, our method constructs a fully differentiable pipeline for jointly optimizing the geometry, appearance, and physical parameters. Our contributions are:
@@
  • A jointly differentiable representation of the shape, appearance, and physical properties of objects.
@@
We evaluate our method on a series of identification problems in both simulated and real-world environments. The results show that our method can infer accurate world models from a single episode of robot interactions with the environment. Finally, we will release the code and data for reproducing the experiments in this paper once the manuscript is accepted.
@@ II-A Differentiable Rigid Body Simulator
Rigid body simulators are essential tools in robotics and engineering for testing, verification, perception, control, and planning. Traditional rigid body simulators are not differentiable, but there have been many recently proposed differentiable rigid body simulators [13, 14, 15, 16, 17, 18] for facilitating downstream system identification, robot planning, and policy optimization tasks. Different strategies are adopted to enable the calculation of gradients for the underlying non-differentiable contact dynamics, including employing a smooth contact model [18, 15, 16], using sub-gradients of the linear complementarity problem [14], and implicit gradients of nonlinear optimization [13, 17]. However, most of these methods do not provide gradients with respect to the geometry, with the exception of  [17, 13, 15]. In this work, we adopt the simulator proposed by Strecke et al. [13] for its physical realism, numerical stability, and fast computation from GPU acceleration.
@@ II-B Learnable Geometry and Appearance Models
Learning 3D geometry and appearance models from 2D raw images is vital to robots’ understanding of the physical world. Earlier research has focused on learning only the 3D geometries without appearance, including point-cloud-based [19] models, convex-hull-based [20] models, and learning implicit signed distance functions [21]. Neural radiance fields (NeRF) is the first method that enables a continuously learnable model for the full 3D appearance of objects and scenes, which uses neural networks to parameterize the spatial appearance properties and implicitly learn the 3D geometry. More recently, Kerbl et al. proposed Gaussian Splatting (GS) [9], a non-parametric method that represents the appearance of the scene with 3D Gaussians and significantly improves the training and rendering speeds due to their fit for fast GPU/CUDA-based rasterization. These algorithms are capable of learning detailed 3D object and scene appearances from sparse image-based observations. However, NeRF and GS lack a clear definition of rigid object surfaces as NeRF represents objects with a neural field and GS with individual 3D Gaussians. Therefore, while there are some initial attempts at integrating them with rigid body simulators [22, 23], research for robust, physically correct, and differentiable collision detections with these models is still ongoing. In addition, NeRF and GS require many diverse views of an object, which is unrealistic in typical robotic manipulation applications.
Compared to standard 3D representations such as point clouds, which do not allow volumetric collision detection, or meshes, which do not allow large geometric and topological changes during optimization, our SaP-based methods enjoy the best of both worlds. Combined with a differentiable renderer, our object representation then achieves end-to-end image-based shape optimization.
@@ II-C World Models
Traditional system identification methods [24] identify only the dynamics parameters from full state information. However, to support diverse downstream robot tasks in the real world, world models need to be built from raw observations and support both accurate dynamics prediction and photorealistic novel view synthesis. While existing works have identified world models from raw image observations using differentiable simulators [17, 13, 15], none supports simultaneous optimization of geometry, appearance, and physical parameters. Another line of work closely related to ours is image-based generative world modeling. These works aim to predict the next RGB frame based on the current frame and action. These models are learned by training on diverse datasets with generative modeling techniques such as variational autoencoders [1, 2] and diffusion [3, 4]. The key differences between our method and these works are that our simulation, grounded in physics, is always physically consistent and is a general-purpose rigid simulator that can provide physical information such as contact forces. Purely data-driven world models generalize poorly to novel scenarios and their lack of physical information severely limits their application to downstream robot tasks. Finally, a recent work [25] proposed a method to use Gaussian Splatting along with a particle-based simulator to track and reconstruct a moving scene. Instead of identifying the physical parameters of the scene, the method optimizes virtual forces attached to each particle such that they match the observed object trajectory. Therefore, although the method can be used as a world model for prediction, the accuracy is severely limited. We include this method as a baseline in our experiments in Sec. V and demonstrate the limitation of this method.
@@
Figure 2: Overview of the proposed fully differentiable pipeline for world model identification from sparse robot observations. Our object representation couples an oriented point cloud $\mathcal{P}$ and a 3D appearance grid $\mathcal{\psi}$. Through a differentiable Poisson solver and differentiable marching cubes, the oriented point cloud is converted to an indicator grid $\chi$ and then a mesh, whose vertex textures are interpolated from the appearance grid $\mathcal{\psi}$. Feeding the object mesh, physical parameters $M$ and $\mu$, the terrain point cloud $\mathcal{P}_{t}$, and the robot pushing trajectory and control $\langle e^{t},u^{t}\rangle$ into a differentiable rigid body simulator and renderer, the predicted scenes can be rendered. Calculating the loss against observed RGB-D images, the scene shape, appearance, and physical parameters are jointly optimized with gradient descent.
@@ III Problem Definition
In this section, we describe our formulation of world model identification. We assume the environment consists of a rigid object and rigid terrain, whose physical properties and appearances are parameterized by $\theta$. A robot, equipped with joint encoders and end-effector force sensors, interacts with the object at $T$ time instances: $t_{1},\cdots,t_{T}$, with a fixed time step $\delta t$. At each time step, the robot observes its end-effector pose $e^{t}\in\text{SE}(3)$ and contact force $f^{t}\in\mathbb{R}^{3}$. Further, the robot is equipped with an RGB-D camera with known intrinsics that observes the object through image $o^{t}\in\mathbb{R}^{H\times W\times 4}$ at camera pose $c^{t}\in\text{SE}(3)$. We further assume an image segmentation mask $m^{t}\in\mathbb{R}^{H\times W\times 4}$ is provided for the object, robot, and terrain. Therefore, the robot observations are a sequence $\mathcal{O}=\{\langle t,e^{t},f^{t},o^{t},c^{t},m^{t}\rangle\}$, and our goal is to estimate $\theta$ from the set of sparse observations $\mathcal{O}$. We formulate this problem as a physics-constrained optimization by introducing a full-fledged physics simulator function $q^{i+1}=g(q^{i},u^{i},\theta)$ that can differentiate through objects’ appearance, geometry, and physical parameters. Here $q^{i}\in\text{SE}(3)\times\text{SE}(3)$ is the object and robot end-effector poses at timestep $t$ and $u^{i}$ is the applied robot force at the end-effector, which is equal in magnitude to the sensed contact force but opposite in direction.
@@ IV-A Differentiable Object Representation
We find that the SaP framework [11], when augmented by additional appearance properties poses an ideal representation for our purpose. Briefly, this framework represents the object using a point cloud with normals on the object surface, denoted as $\mathcal{P}=\{(p\in\mathbb{R}^{3},n\in\mathbb{R}^{3})\}$. These normal directions induce a discrete vector field $v(x)=\sum_{\langle p,n\rangle\in\mathcal{P}}n\mathbb{I}[x=p]$. SaP then uses Poisson reconstruction [26] to recover an underlying implicit indicator field $\chi(x)$ that describes the occupancy of the solid geometry, i.e. whether $x$ is inside or outside the geometry, and matches its gradient field with $v(x)$ by solving the variational problem:
@@
which amounts to solving the Poisson equation $\Delta\chi=\nabla\dot{v}$. SaP discretizes the indicator field $\chi$ on a uniform grid domain $\Omega$, which allows the efficient solution of $\chi$ via GPU-accelerated Fast Fourier Transform (FFT) with well-defined derivatives. We use a $128\times 128\times 128$ discretized grid $\chi$ for all the experiments in this paper.
The indicator field $\chi$ is then transformed to a triangle mesh $\mathcal{M}$ with a differentiable marching cubes algorithm [12]. Collision detection can then be easily achieved with standard techniques for meshes. To enable appearance modeling, we further augment with a grid of appearance properties, with the same grid resolution as the one storing the indicator field $\chi$. The appearance property is then propagated to the mesh vertices via tri-linear interpolation. In this work, we only store and render the color field, denoted as $\psi$, but other appearance properties can be incorporated in the same manner as required by more advanced differentiable rendering equations. The mesh can then be rendered using any differentiable renderer framework such as [27, 28], for which we use the open source implementation in PyTorch3D [29]. Specifically, at the time instance $t$, we invoke the renderer with the object transformed to $q^{t}$ and the camera transformed to $c^{t}$. Our parametrization of the object’s physics and appearance is defined as:
@@
where the first two parameters are mass-inertial properties and frictional coefficients, defining the physical properties, and the last two parameters are the oriented point cloud for SaP and color field, defining the objects’ appearance.
@@ IV-B Differentiable Simulator
For our application, we only consider unconstrained rigid body dynamics with dry frictional contacts. Note that additional physical constraints for describing objects such as soft bodies and articulated objects can be potentially incorporated into our framework and its differentiation has been well-studied, e.g. in [16].
@@
with $M(q^{i})$ being the generalized mass matrix, $C(q^{i},\dot{q}^{i})$ being the centrifugal, Coriolis, and gravitational force, $J^{i},J^{\perp},J^{\parallel}$ being the Jacobian matrix for the external, normal, and tangent contact forces at all the detected contact points, respectively. Finally, $\tau^{\perp},\tau^{\parallel}$ are the contact forces. At each time step, a mixed linear complementarity problem (LCP) is solved to calculate the constraint forces $\tau^{\perp},\tau^{\parallel}$, yielding the final acceleration $\ddot{q}^{i}$, and we then integrate the configuration forward in time [32, 33] as:
@@
$\begin{cases}0\leq\tau^{\perp}&\perp J^{\perp T}\dot{q}^{i+1}\geq 0\\ 0\leq\tau^{\parallel}&\perp\lambda e+J^{\parallel T}\dot{q}^{i+1}\geq 0\\ 0\leq\lambda&\perp\mu\tau^{\perp}-e^{T}\tau^{\parallel}\geq 0,\end{cases}$ (4)
with $e$ being the unit vector and $\mu$ being the frictional coefficient. $\lambda$ is an auxiliary variable encoding the stick or slip frictional state. To differentiate through the simulator, we adopt the differentiation technique proposed by [14, 13], where the result of the LCP is made differentiable by solving with a primal-dual method and performing sensitivity analysis at the solution to yield derivatives with respect to the problem data. In this way, the derivatives propagate the gradient information to the Jacobian matrix $J^{\perp,\parallel}$, and finally to the object geometric parameters $\mathcal{P}$. In summary, Eqn. 2,3,4 defines our differentiable simulator function $g$. In particular, we adopt the differentiable simulator proposed by Strecke et al. [13] for its fast implementation on GPU.
Simulation Real World
Figure 3: The experiment setups for the simulated (left) and physical (right) experiments. 9 objects are used for simulation with the PyBullet simulator, including 8 YCB objects and a green box. For the real-world experiments, three YCB objects (Drill, Mustard, and Sugar) are used. A UR5e arm equipped with a pusher and an overhead Realsense D435 RGB-D camera are used. Note that only the circled object in the real-world setup is the object of interest and everything else is treated as the static terrain.